# Forest-Chat: Adapting Vision-Language Agents for Interactive Forest Change Analysis

### v1

@@ 5.3 Forest-Chat Zero-shot Performance
The remaining hyperparameters are not tuned due to their minimal impact on mask generation. Points-per-side - which controls the spatial density of point prompts provided to SAM for mask proposal generation - is fixed at 16 as a practical trade-off between performance and run-time. Performing a full hyperparameter sweep on LEVIR-MCI-Trees is computationally prohibitive, so the best configuration from Forest-Change is reused. These parameters mainly control proposal density and filtering heuristics, rather than dataset-specific semantics, so reusing them is unlikely to introduce systematic bias.
Table 3 provides an overview of FC-Zero-shot’s change detection performance. Of the hyperparameters, the change confidence threshold was found to have the highest influence on performance. Overall, the optimal hyperparameters differed minimally from those found by Zheng et al. (2024), with performance increases of 0.95% MIoU and 9.28% on the Forest-Change and LEVIR-MCI-Trees datasets respectively. Zero-shot performance is noticeably lower on both datasets when compared to the supervised methods, but the discrepancy is much more pronounced for LEVIR-MCI-Trees. This can be partly attributed to the segmentation masks only containing annotations for road and building changes, so any other identified changes will be penalised during evaluation. Additionally, FC-Supervised has a stronger supervision signal on the LEVIR-MCI-Trees dataset due to better image quality and variety, resulting in higher segmentation quality. A visual comparison between the various benchmarked methods on the Forest-Change dataset is provided in Figure 6.
Generally, FC-Zero-shot is highly sensitive to atmospheric artifacts and seasonal change, resulting in frequent cases of false positives. This sensitivity can potentially be mitigated by providing point queries. However, point queries are best utilised when segmenting over multiple object classes, as compared to the binary case presented by Forest-Change. Besides, there is no methodological process for determining the optimal placement of point queries. However, the development of an interactive interface that enables users to provide point queries is a useful contribution for faster investigations into optimal point prompt placements.
Table 3: FC-Zero-shot segmentation performance with default parameters and best hyperparameter sweep configuration. Best configuration: change confidence threshold = 145; stability score threshold = 0.93; area threshold = 0.9; object similarity threshold = 60.
Dataset Model MIoU $IoU_{\textit{nc}}$ $IoU_{\textit{c}}$
Forest-Change FC-Zero-shotD 58.56 93.83 23.29
FC-Zero-shotB 59.51 94.64 24.38
FC-Supervised 67.10 96.12 38.07
LEVIR-MCI-Trees FC-Zero-shotD 38.04 54.05 22.04
FC-Zero-shotB 47.32 66.60 28.04
FC-Supervised 88.13 95.89 80.36
@@
Despite lowering the points-per-side hyperparameter, inference time is still significantly slower than all other benchmarked methods, as shown in Table 2. FC-Zero-shot runs the SAM image encoder on each image, using a ViT_h as the default backbone, resulting in extensive inference times. Reducing the ViT size would speed up inference, but compromise performance, which is already a limiting factor to practical applications. While this approach is capable of outperforming pre-VLM supervised change detection approaches (e.g. (Hewarathna et al., 2024)), and has the ability to detect changes that may have been missed by human annotators, it lags behind contemporary supervised methods and is sensitive to atmospheric noise and seasonal variations. Future work in this area could focus on extending SAM’s (Kirillov et al., 2023) textual prompting to change detection, or accounting for challenges inherent to RS imagery, something that the underlying SAM model is not designed for due to pre-training focusing on natural images. Comparing alternative remote sensing-pretrained foundational vision backbones would be valuable for aligning extracted features with change mask proposals.
5.4 Balancing Change Detection and Captioning Study
@@
While normalised loss balancing is effective at preventing one task from dominating the learning gradient, it statically assumes that tasks are equally important and difficult throughout training, potentially ignoring the fundamental noise heteroscedasticity between tasks. To investigate this, we experiment with various multi-task learning (MTL) loss balancing and gradient normalisation techniques to assess their impact on the joint training of CDC and their potential to improve collaborative learning.
@@
For MTL loss balancing, we select Uncertainty Weighting (Kendall et al., 2018) and Enhanced Dynamic Weight Averaging (EDWA) (Shi et al., 2024). For gradient conflict resolution, we evaluate Conflict-Averse Gradient Descent (CAGrad) (Liu et al., 2021), Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020), and GradDrop (Chen et al., 2020). These methods are also combined in various configurations to form a comprehensive experimental suite capable of detecting any combinations that produce notable performance differences.
Experiments are performed only on the smaller Forest-Change dataset due to computational limitations. Due to the stochastic nature of the full experimental results presented in Tables 1 and 2, the results for each method are aggregated and presented in Table 4. The results show clearly that the optimum strategy previously found of normalising the losses against each other remains the best approach for balancing the two tasks. This suggests that the FC-Supervised’s learning objectives are naturally aligned, and that the architecture already provides sufficient task decoupling. Although the captioning loss exhibits larger numerical values than the change detection loss, equal task weighting and no gradient conflict resolution generally yields superior performance. This suggests that loss magnitude alone is not indicative of effective gradient influence. In practice, the dense spatial supervision of change detection produces strong and stable gradients that naturally balance the higher-variance sequence loss from captioning. In these circumstances, dynamic loss balancing and gradient conflict resolution largely introduce unnecessary optimisation noise, potentially disrupting beneficial cross-task regularisation. Uncertainty weighting is the only approach that marginally improves some captioning metrics, but at a cost to segmentation. These results validate the efficacy of a joint learning framework for RSCDC, with future work likely to benefit from designing better model architectures and training strategies enhanced task performance.
@@
Table 4: Aggregated change detection and change captioning performance for MTL strategies on the Forest-Change dataset. Mean $\pm$ standard deviation across configurations, with three runs each. Bold indicates the best mean value per metric within each group when the improvement is clearly distinguishable relative to standard deviation
@@
MTL Strategy Method MIoU $IoU_{\textit{nc}}$ $IoU_{\textit{c}}$ B1 B2 B3 B4 METEOR ROUGEL CIDEr-D
@@
5.5 FC-Supervised Backbone Size Influence
This section reports on the importance of the Segformer backbone model for feature extraction on downstream tasks. By default, a MiT-B1 is used, but this study tests change detection and captioning performance when using the B0 and B2 sizes. Due to computational limitations and the diminishing returns observed in the original Segformer paper (Xie et al., 2021), backbone sizes above B2 are not explored in this study. Table 5 clearly displays a trend of increasing performance in change detection for both datasets as the backbone size increases. FC-Supervised with the MiT-B2 backbone is shown to be the best change detection model across both datasets, with greater backbone sizes driving improved change class identification. On the Forest-Change dataset, the Mit-B2 backbone outperforms BiFA (Zhang et al., 2024a) in Table 1, making FC-Supervised the best performing model when using a larger backbone. Beyond Mit-B1 though, overall performance gains begin to diminish on both tasks, and even recede for change captioning, indicating that the features extracted by larger backbone predominantly benefit the change detection task more. Training times also grow exponentially, limiting the usage of backbones beyond MiT-B2. Despite slower inference times, FC-Supervised with the MiT-B2 backbone trains at a speed comparable to the benchmark models from Section 5.2, highlighting the efficiency of the architecture’s dual-training approach. These results indicate that backbones need to be carefully selected based on the ability to extract features equally useful to both tasks - especially from models with fewer parameters to improve practical usage.
Table 5: Change Detection and Change Captioning performances on the test sets of LEVIR-MCI-Trees and Forest-Change datasets using different Segformer backbone sizes for FC-Supervised. Best results per metric are bold, second-best are underlined. The training time is reported as a rounded mean $\pm$ standard deviation (s) for a single epoch. Results are averages of three runs.
@@ 5.5 FC-Supervised Backbone Size Influence
5.6 Forest-Chat as a Tool for Forest Change Analysis

### v2

@@ 5.3 Forest-Chat Zero-shot Change Detection
The resulting segmentation performance for each backbone and configuration is summarised in Table 3. Among the evaluated hyperparameters, the change confidence threshold has the greatest influence on performance, as it directly controls the sensitivity of the cosine-distance based change detection mechanism. Across the evaluated backbones, hyperparameter tuning consistently improves segmentation quality across both classes by increasing sensitivity to genuine change regions while suppressing spurious detections caused by atmospheric artifacts and seasonal variations.
Zero-shot performance remains substantially lower than the supervised FC-Supervised baseline across both datasets, with the gap particularly pronounced on the LEVIR-MCI-Trees dataset. This discrepancy can partly be attributed to the dataset’s annotation protocol, which only labels road and building changes. Consequently, other legitimate changes detected by the model are penalised during evaluation, contributing to the dataset’s observed sensitivity to hyperparameter selection. In contrast, the Forest-Change dataset adopts a more consistent definition of change, resulting in comparatively stronger zero-shot performance and greater robustness to hyperparameter variation.
Table 3: FC-Zero-shot segmentation performance with default parameters and best hyperparameter sweep configuration for the SAM (ViT-B and ViT-H) and SAM2 (SAM 2.1 Hiera Large) models. SAM Best configuration : change confidence threshold = 145; stability score threshold = 0.93; area threshold = 0.9; object similarity threshold = 60. SAM2 Best configuration : change confidence threshold = 150; stability score threshold = 0.91; area threshold = 0.9; object similarity threshold = 50. P = precision, R = recall.
Dataset Model mIoU $IoU_{\textit{nc}}$ $IoU_{\textit{c}}$ P R
[…]
FC-Supervised 88.13 95.89 80.36 94.30 92.92
[…]
FC-Supervised 67.10 96.12 38.07 84.19 76.82
@@
Backbone selection introduces a clear trade-off between segmentation quality, detection sensitivity, and computational cost. As shown in Table 3, the larger ViT-H backbone generally achieves stronger $IoU_{\textit{c}}$ performance, indicating improved sensitivity to subtle or fragmented change regions. However, this increased sensitivity can also amplify false detections caused by illumination shifts, cloud cover, or phenological differences between image acquisition dates, as illustrated qualitatively in Figure 7. This behaviour is reflected in the precision–recall characteristics reported in Table 3. On the LEVIR-MCI-Trees dataset, the best-performing ViT-H configuration achieves higher recall (70.34%) than ViT-B (61.49%), indicating greater sensitivity to change regions, but with only marginal improvements in precision. A similar trend is observed on the Forest-Change dataset, where ViT-H attains higher recall (66.69%) but lower precision (67.97%) relative to the more conservative ViT-B configuration, which achieves higher precision (80.93%) at the cost of reduced recall (62.69%).
The SAM2 backbone provides an intermediate alternative, achieving comparable overall performance while partially reducing the computational burden relative to larger encoders such as ViT-H. However, $IoU_{\textit{c}}$ performance remains comparatively limited. Although SAM2 introduces architectural improvements designed to improve efficiency and temporal consistency in video segmentation, these advantages are not fully utilised in the present pipeline, which processes the two temporal images independently and relies primarily on feature similarity for change inference. Consequently, the SAM2 configuration does not consistently outperform the original Segment Anything Model backbones in this setting.
Despite lowering the points-per-side hyperparameter, inference time remains substantially slower than other benchmarked methods (Table 2), limiting practical deployment in large-scale monitoring scenarios. While smaller backbones can partially reduce this computational burden, the associated decline in segmentation quality limits the practical benefits of this trade-off. In practice, improvements in segmentation sensitivity do not always correspond to cleaner change maps, particularly when radiometric, atmospheric, or seasonal variations are interpreted as genuine change.
These results suggest that purely zero-shot approaches may have inherent limitations for reliable change detection in remote sensing imagery. Rather than relying exclusively on zero-shot inference, future work may benefit from exploring architectures that combine strong pretrained encoders with lightweight adaptation mechanisms capable of learning from limited labelled data. Approaches such as META-CD Gao et al. (2025) demonstrate that applying domain-adaptive pretraining for vision foundation models can significantly improve change detection performance while maintaining strong generalisation under limited supervision, and lighter backbones could be selected to balance performance against inference times.
5.4 Forest-Chat Zero-shot for Change Captioning
and Refinement
Table 4: FC-Zero-shot Change Captioning and Caption Refinement on the test sets of LEVIR-MCI-Trees and Forest-Change datasets. Results for refined captions are an average of three runs.
Dataset Model B1 B2 B3 B4 METEOR ROUGEL CIDEr-D BERTScore
LEVIR-MCI-Trees FC-Zero-shotG 52.29 31.90 17.89 10.19 18.00 32.94 18.86 86.42
FC-Zero-shotS 67.56 45.68 29.45 18.23 21.19 37.20 29.95 84.27
FC-Supervised 75.25 60.90 46.21 34.41 23.32 49.34 48.69 86.34
FC-Supervised-Refined 59.49 40.57 27.14 17.54 22.90 38.81 18.49 87.36
Forest-Change FC-Zero-shotG 25.88 10.86 3.65 0.00 8.99 21.11 2.06 82.82
FC-Zero-shotS 52.97 43.64 38.20 34.00 24.96 39.87 32.13 82.47
FC-Supervised 67.54 56.34 47.55 40.17 28.22 48.52 38.79 85.63
FC-Supervised-Refined 73.19 59.87 50.63 43.31 32.98 54.19 50.83 86.73
G General prompt for FC-Zero-shot;  S Dataset-specific style-guided prompt for FC-Zero-shot.
Two prompting strategies are compared: a general prompt providing minimal task context, and a dataset-specific style-guided prompt designed to align FC-Zero-shot’s outputs with the vocabulary, structure, and thematic content of each dataset. As shown in Table 4, style-guided prompting substantially improves zero-shot performance across both datasets, with BLEU-4 increasing from 10.19 to 18.23 on LEVIR-MCI-Trees and from 0.00 to 34.00 on Forest-Change. The latter result highlights that zero-shot general prompting alone is insufficient to produce captions consistent with the semantic and stylistic conventions of the target dataset. Despite these gains, FC-Zero-shot under style-guided prompting remains below FC-Supervised on most metrics, confirming that zero-shot MLLMs cannot yet supplant supervised change captioning models in specialised remote sensing domains.
While zero-shot performance is limited, FC-Zero-shot can still provide value as a targeted refinement stage for FC-Supervised outputs, with the extent of benefit varying between datasets. On Forest-Change, refinement consistently improves captions. The reference captions are largely rule-based, which causes much of the semantic detail available in the human-written captions to be overlooked. Refinement enriches the captions with additional spatial and contextual information, which is reflected in improvements in BLEU-4 (40.17 → 43.31) and CIDEr-D (38.79 → 50.83), while preserving the underlying change semantics. In contrast, refinement on LEVIR-MCI-Trees introduces additional descriptive detail that, while semantically accurate, diverges lexically from the concise human-written references. This results in declines in n-gram metrics (BLEU-4: 34.41 -¿ 17.54, CIDEr-D: 48.69 -¿ 18.49) while BERTScore improves slightly (86.34 → 87.36). This demonstrates that semantic content is preserved even when lexical overlap decreases. These results indicate that refinement is practically valuable as a targeted post-processing step for known supervised model limitations (e.g. correcting geographic positioning, fixing grammatical inconsistencies, or incorporating domain-specific vocabulary), rather than as a general-purpose quality enhancement. Tables 1 and 2 in the appendix present selected examples illustrating the spatial and semantic enrichment introduced by FC-Zero-shot refinement over FC-Supervised predicted captions on the Forest-Change and LEVIR-MCI-Trees test sets respectively.
Overall, these results show that multimodal LLMs can be effective for change captioning, particularly when guided by dataset-specific prompts or used for targeted refinement. Zero-shot utility remains limited without explicit domain grounding, and the benefits of refinement are dependent on the reference caption characteristics. Forest-Chat’s innate supervised captioning remains strong, but the addition of zero-shot captioning and refinement offers significant utility to users. Future work should broaden benchmarking of open-source models within the domain, and investigate fine-tuning strategies or incorporating domain-aware knowledge bases to better align LLM outputs with the structural and vocabulary constraints of remote sensing change captioning datasets.
5.5 Cross-Domain Adaptation Study
Assessing model generalisation across diverse remote sensing scenes and sensors is critical for understanding downstream performance and real-world operational behaviour. Consequently, this section evaluates the change detection models benchmarked in Section 5.2 by first pretraining on Forest-Change and then fine-tuning on varying proportions of the JL1-CD-Trees training set, from 5% up to 100%, before evaluation on its test set. FC-Zero-shot serves as a training-free baseline. For FC-Supervised, we additionally compare pretraining on Forest-Change using either full multi-task learning or change-detection-only to assess the transferability of task-specific versus multi-task features. We also evaluate each model when trained on the full JL1-CD-Trees dataset with and without Forest-Change pretraining, to directly quantify how effectively pretrained features improve performance in the new domain.
Figure 6: Cross-domain transfer performance on JL1-CD-Trees dataset. a) Models pretrained on Forest-Change and fine-tuned on increasing fractions of JL1-CD-Trees. b) Comparison of models trained solely on JL1-CD-Trees versus pretrained on Forest-Change and fine-tuned on the full train set. FC-Zero-shot is included as a baseline.
The performance trends observed on the JL1-CD-Trees dataset in Figure 6 reveal clear patterns in how the benchmarked architectures adapt to cross-resolution target-domain data. Zero-shot transfer from Forest-Change pretraining yields consistently poor performance across all supervised models, with a particularly severe collapse for U-Net SiamDiff. In contrast, FC-Zero-shot achieves an mIoU of 54.98 using Forest-Change-optimised hyperparameters, demonstrating consistent cross-domain change detection capabilities under training-free conditions. The near-uniform collapse in zero-shot performance across supervised models underscores the severity of the domain gap between Forest-Change and JL1-CD-Trees, driven most consequentially by the order-of-magnitude difference in spatial resolution (30 m vs ¡1 m) and the absence of matched seasonal representation in the source domain.
However, once target-domain supervision is introduced, performance improves rapidly across all supervised architectures. Even with only 5–10% of the JL1-CD-Trees training data, mIoU rises substantially from zero-shot levels, where most models initially default to predicting no change, or behave near-randomly as in the case of U-Net SiamDiff. This indicates that while pretrained features provide a useful starting point, models require even modest target-domain supervision to correctly interpret high-resolution change patterns. Performance continues to increase steadily as more data is added, with most architectures approaching mIoU values in the low-to-mid 70s when trained on the full dataset. Every model exhibits a notable performance jump between 10% and 25% of the training data, after which improvements become more incremental, suggesting that a relatively small amount of target-domain supervision is sufficient to recalibrate models to the spatial characteristics of the JL1-CD-Trees imagery, with additional data primarily refining boundary precision and small-object detection.
The role of Forest-Change pretraining is more nuanced. While pretraining accelerates early-stage adaptation, it does not consistently improve final performance when the full JL1-CD-Trees training set is available. In several cases, models trained from scratch on JL1-CD-Trees achieve comparable or slightly higher performance than their pretrained counterparts. FC-Supervised provides a notable example, achieving stronger results when trained directly on JL1-CD-Trees than when initialised with Forest-Change pretraining, suggesting that coarse-resolution representations may introduce biases that hinder adaptation to the finer spatial structures present in the target dataset. In contrast, architectures such as BiFA and Change3D benefit more consistently from increasing supervision when initialised with Forest-Change pretraining, with BiFA achieving the strongest performance among the pretrained models when trained on the full dataset. Differences between the multi-task and CD-only FC-Supervised variants are comparatively small, indicating that the additional captioning supervision does not substantially alter transfer behaviour for the detection task. Overall, these findings suggest that while pretraining on related RS datasets can provide useful initialisation, its effectiveness depends strongly on the quality of the source dataset, and the spatial alignment between source and target domains.
5.6 Ablation Studies
5.6.1 Multi-task Loss Balancing and Gradient Conflict
Resolution Study
@@ 5.6.1 Multi-task Loss Balancing and Gradient Conflict
As discussed in Section 4.5, joint optimisation of change detection and captioning may introduce gradient interference due to the differing objectives of spatial localisation and semantic language generation. While normalised loss balancing is effective at preventing one task from dominating the learning gradient, it assumes that tasks are equally important and difficult throughout training, potentially ignoring the fundamental noise heteroscedasticity between tasks. To investigate this, we experiment with various multi-task learning (MTL) loss balancing and gradient conflict resolution techniques to assess their impact on the joint training of CDC and their potential to improve collaborative learning. Several commonly used gradient adjustment methods are compared to examine their influence on multi-task optimisation.
@@
For MTL loss balancing, we select Uncertainty Weighting (Kendall et al., 2018) and Enhanced Dynamic Weight Averaging (EDWA) (Shi et al., 2024). For gradient conflict resolution, we evaluate Conflict-Averse Gradient Descent (CAGrad) (Liu et al., 2021), Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020), and GradDrop (Chen et al., 2020). To ensure a comprehensive experiment suite, every methodological combination is explored.
@@
Table 5: Aggregated change detection and change captioning performance for MTL / Gradient Conflict resolution strategies. Mean $\pm$ standard deviation across configurations, with three runs each. Bold indicates the best mean value per dataset per metric within each method group. When results are tied or close, preference is given to the method with the lower standard deviation.
@@
Dataset MTL / Gradient Conflict Strategy mIoU $IoU_{\textit{nc}}$ $IoU_{\textit{c}}$ B1 B2 B3 B4 METEOR ROUGEL CIDEr-D
@@
[…]
Experiments are performed across the Forest-Change and LEVIR-MCI-Trees datasets to assess sensitivity to data domain. Due to the stochastic nature of the full experimental results presented in Tables  14, the results for each method are aggregated and presented in Table 5. Across both datasets, segmentation performance remains largely stable regardless of the multi-task learning optimisation strategy, with Equal weighting only having an mIoU improvement of ~0.3-0.5 over Uncertainty. In contrast, captioning metrics exhibit greater variability, indicating that multi-task interactions primarily affect the language generation component. Here, Uncertainty is generally the better-performing MTL loss balancing method, with BLEU-4 ~0.7 greater than Equal losses. For the majority of metrics, no gradient conflict resolution techniques are required to stabilise task harmony. The only exception to this is Table 2, in which Uncertainty with CAGrad on the LEVIR-MCI-Trees dataset displays notable improvements for both tasks, but this result is not replicated on Forest-Change. The observed invariance of change detection performance and the sensitivity of change captioning to MTL loss balancing strategy is likely due to the dense supervision and complementary nature of the segmentation task with FC-Supervised, which stabilises shared feature representations. Conversely, the captioning task receives sparser, sequential supervision and is more sensitive to variations in gradient contributions. These results suggest that, in settings where tasks are naturally aligned, careful architectural design and shared representations may be more important than sophisticated multi-task optimisation strategies. In such cases, loss balancing and gradient conflict resolution can introduce unnecessary optimisation noise, potentially disrupting beneficial cross-task regularisation. Developing architectures that allow tasks to reinforce each other during joint learning should therefore be prioritised over relying on complex optimisation techniques. Future work could also explore meta-learning based task weighting approaches, such as Auto-Lambda (Liu et al., 2022b), which dynamically learn task weights through meta-optimisation and may further improve multi-task synergy.
5.6.2 Forest Chat Supervised Backbone Size Influence
This section reports on the importance of the Segformer backbone model for feature extraction on downstream tasks. By default, a MiT-B1 is used, but this study evaluates change detection and captioning performance using the smaller MiT-B0 and larger MiT-B2 variants. To provide a lightweight convolutional comparison, a RegNetX-400MF (Radosavovic et al., 2020) backbone is also evaluated. Due to computational limitations and the diminishing returns observed in the original SegFormer paper (Xie et al., 2021), backbone sizes above B2 are not explored in this study. Table 6 displays a general trend of increasing change detection performance for both the LEVIR-MCI-Trees and Forest-Change datasets. FC-Supervised with the MiT-B2 backbone achieves the strongest change detection performance across both datasets, with larger backbone sizes improving change class identification. On the Forest-Change dataset, the MiT-B2 configuration also outperforms BiFA (Zhang et al., 2024a) in Table 1, making FC-Supervised the best performing model when using a larger backbone.
Table 6: Change Detection and Change Captioning performances on the test sets of LEVIR-MCI-Trees and Forest-Change datasets using different Segformer (MiT-Bn) backbone sizes and RegNetX-400MF for FC-Supervised. Best results per metric are bold, second-best are underlined. The training time is reported as a rounded mean $\pm$ standard deviation (s) for a single epoch. Results are averages of three runs.
@@ 5.6.2 Forest Chat Supervised Backbone Size Influence
However, performance gains begin to diminish beyond the MiT-B1 backbone. Improvements from MiT-B1 to MiT-B2 are relatively small, and captioning metrics show little benefit from the larger backbone, suggesting that the additional representational capacity predominantly benefits the change detection task. The lightweight RegNetX-400MF backbone produces the lowest change detection performance, especially for change class discrimination, while demonstrating notably strong captioning performance. On LEVIR-MCI-Trees, RegNetX-400MF outperforms MiT-B0 across most metrics, while on Forest-Change it achieves some of the strongest captioning scores across several metrics, including higher BLEU-1 and CIDEr-D scores than the MiT-B2 backbone. This suggests that the semantic information required for caption generation can still be captured effectively with a compact convolutional feature extractor when supported by the cross-modal reasoning components of the architecture. In contrast, improvements in backbone capacity primarily benefit the spatial localisation required for accurate change detection.
Training times also grow exponentially as backbone size increases, limiting the practicality of configurations beyond MiT-B2. While larger backbones offer modest detection improvements, the gains diminish relative to the associated computational cost. Despite slower inference times, FC-Supervised with the MiT-B2 backbone trains at a speed comparable to the benchmark models from Section 5.2, highlighting the efficiency of the architecture’s dual-training approach. Overall, these results indicate that backbone selection should balance representational capacity and computational efficiency, prioritising architectures that extract features useful to both change detection and captioning tasks while maintaining practical training requirements.
5.6.3 Forest-Chat as a Tool for Forest Change Analysis